# Benchmarking Filtered Approximate Nearest Neighbor Search Algorithms on Transformer-based Embedding Vectors

### v1

@@ 3.1. Filtering Approaches
The efficiency of different approaches to solving the FANNS problem depends on the dataset and filtering condition (see Figure 2). We define the selectivity of a filter $f$ on a dataset $D=\{I_{i}\mid i\in\{1,\dots,n\}\}$ as in previous work [^1] Note that some works (Wei et al., 2020; Wang et al., 2021d; Xu et al., 2024) use the inverse definition of selectivity. (Mohoney et al., 2023; Zhang et al., 2023; Patel et al., 2024; Liang et al., 2024):
$\text{selectivity}=\frac{|\{I_{i}\mid I_{i}\in D\land f(I_{i})=\texttt{true}\}|}{|D|}.$
In other words, selectivity is the fraction of database items that match the filter. Pre-filtering is most efficient when selectivity is low, meaning only a few items remain for the ENNS. Post-filtering and in-filtering are most efficient when selectivity is high, i.e., most items pass the filter. If selectivity is too low, ANNS may not return enough valid candidates for post-filtering, and graph traversal with in-filtering may fail due to a sparsely connected or disconnected graph. A hybrid index is most efficient when selectivity is moderate, making none of the three previous approaches optimal.
3.2. Classification of FANNS Methods
We classify FANNS methods along three dimensions:
  • Approach: Each FANNS method follows one or a combination of the following approaches: pre-filtering, post-filtering, in-filtering, or hybrid indexing (see Section 3.1).
  • Technique: FANNS methods build upon the same four fundamental techniques as ANNS methods: tree-based, hash-based, graph-based, or quantization-based (see Section 2.5).
  • Filter Type: Different FANNS methods support various filter types (see Section 2.4 for details).
Table 3 provides an overview of existing FANNS methods classified along these three dimensions.
Figure 2. 3.1) The optimal approach for solving FANNS depends on the filter’s selectivity within a given dataset.
3.3. Explanation of FANNS Methods
@@ 3.3.5. Milvus
Milvus (Wang et al., 2021d) extends the FAISS ANNS library (Douze et al., 2024; Research, 2025) by incorporating range filtering for one or multiple attributes. It supports both pre- and post-filtering, as well as a partition-based filtering scheme that partitions the data based on the most frequently filtered attribute and builds a separate ANNS index for each partition.
@@ 3.3.6. NHQ
NHQ (Wang et al., 2024) proposes a method to extend any proximity graph-based ANNS index into a hybrid FANNS index. Instead of constructing the ANNS index solely based on embedding distance, it introduces a fusion distance that combines embedding distance with attribute dissimilarity. While this approach is applicable to any proximity graph-based ANNS index, NHQ also presents construction schemes for two novel proximity graphs. Since its query execution scheme operates using the fusion distance without an explicit filtering step, NHQ may return items that do not match the filter. Thus, it addresses the AFANNS problem rather than the FANNS problem.
@@ 3.3.7. HQANN
Like NHQ (Wang et al., 2024), HQANN (Yang et al., 2020c) employs a fusion distance to transform any proximity graph-based ANNS index into a hybrid FANNS index. The key difference between HQANN and NHQ is that HQANN prioritizes attribute dissimilarity in the fusion distance, whereas NHQ emphasizes embedding distance as the dominant factor.
@@ 3.3.8. AIRSHIP
AIRSHIP (Zhao et al., 2022) (see Figure 3) is a graph-based FANNS index capable of handling arbitrary filter functions. It relies on a graph-based ANNS index constructed without considering attributes but adapts query execution to FANNS through in-filtering and a series of optimizations. During query execution, AIRSHIP traverses the entire graph but includes only vertices that satisfy the filter in the result set. Additionally, it selects a starting point within a cluster of items guaranteed to match the filter and traverses the graph concurrently in multiple directions.
@@ 3.3.9. HQI
Apple’s HQI (Mohoney et al., 2023) (see Figure 4) is a hybrid, workload-aware index designed for batch processing of FANNS queries. To integrate attribute filtering with nearest neighbor search, HQI first transforms embedding vectors into attributes by applying k-means clustering and storing each item’s cluster ID as an attribute. It then employs a balanced qd-tree (Yang et al., 2020a), an attribute-only index, to partition the items. Each non-leaf node in the qd-tree contains a predicate over the item attributes (including the cluster ID), with its two child nodes holding items that do or do not satisfy the predicate. The tree’s leaves form a non-overlapping partition of all items. During query execution, HQI identifies the query vector’s $w$ closest cluster IDs and prunes all branches of the tree that do not contain any of these $w$ cluster IDs or whose attributes do not match the filter. An IVF index is used within each leaf.
@@ 3.3.10. VBASE
VBASE (Zhang et al., 2023) extends PostgreSQL (Group, 2025) by integrating ANNS into the database system. Its key insight is that most ANNS query execution methods follow a two-stage process: an initial stage where traversal moves from a starting point in the direction of the query vector, and a second stage where the search gradually moves away from the query vector to identify its KNN. Rather than employing a naïve pre-filtering approach, where the ANNS index is queried with a fixed $k^{\prime}\geq k$ and non-matching items are filtered out, VBASE refines query execution by running the ANNS query only until the traversal starts moving away from the query vector. From that point onward, it iteratively retrieves and filters additional items until the required $k$ matches are found.
@@ 3.3.11. FDANN
Microsoft’s Filtered-DiskANN (FDANN) (Gollapudi et al., 2023) (see Figure 5) implements a graph-based index with EMIS filtering. In addition to supporting an EMIS filter with a single query-label, it also allows filtering with multiple query-labels, where at least one of them must be present in the item’s set attribute. Filtered-DiskANN introduces two methods for constructing FANNS indices based on Vamana graphs (Jayaram Subramanya et al., 2019): FilteredVamana and StitchedVamana. During query execution, it traverses the graph starting from a set of entry points that are guaranteed to contain the query-label. Throughout its NSW-like graph traversal, it considers only those vertices that satisfy the filter criteria.
@@ 3.3.12. CAPS
In contrast to most FANNS methods, CAPS (Gupta et al., 2023) (see Figure 6) introduces a quantization- and tree-based index that is easier to parallelize than the more common graph-based indices. The CAPS index consists of two levels. The first level organizes items into clusters, for example, using an IVF (Sivic and Zisserman, 2003b). This clustering can be based solely on embedding distance or on a combination of embedding distance and attribute dissimilarity. At the second level, CAPS constructs an attribute frequency tree (AFT), a tree-based attribute-only index. Each level of the AFT splits items into two buckets: one containing items with the most common attribute value (a leaf node) and another with the remaining items (a non-leaf node), unless the maximum depth is reached.
@@
During query execution, CAPS first selects the $w$ clusters closest to the query vector (first level). Within these $w$ clusters, it uses AFTs to identify items matching the filter and then performs ENNS on the filtered results (second level).
@@ 3.3.13. ARKGraph
[…]
@@ 3.3.14. ACORN
ACORN (Patel et al., 2024) (see Figure 8) supports arbitrary filter types and filtering across multiple attributes. The ACORN index is a denser variant of the HNSW ANNS index (Malkov and Yashunin, 2018), constructed without considering attributes. During query execution, ACORN traverses only the subgraph induced by vertices that match the filter. This approach differs from AIRSHIP (Zhao et al., 2022), which traverses the entire graph but includes only matching vertices in the result set. The ACORN index is designed such that its subgraphs approximate an HNSW index (Malkov and Yashunin, 2018), ensuring efficient nearest neighbor search while maintaining filtering constraints.
@@ 3.3.15. $\beta$-WST
$\beta$-WST (Engels et al., 2024) (see Figure 9) is a FANNS index designed for range filtering on a single attribute. A $\beta$-WST index is structured as a segment tree (De Berg, 2000) with a uniform branching factor of $\beta$ (see Figure 9). At layer $l$, the attribute value range is divided into $\beta^{l}$ segments, with a graph-based ANNS index constructed for each segment. The authors propose multiple query strategies for $\beta$-WST. The default strategy selects a minimal set of tree nodes that fully cover the query range, queries the corresponding ANNS indices, and merges the results. Additional strategies include OptimizedPostfiltering, TreeSplit, and SuperPostfiltering.
@@ 3.3.16. SeRF
SeRF (Zuo et al., 2024) introduces the segment graph, a data structure designed to efficiently answer range-filtered FANNS queries with half-bounded query ranges (i.e., where the query range has an upper limit but no lower limit). The segment graph losslessly compresses $n$ HNSW indices (Malkov and Yashunin, 2018), corresponding to the $n$ possible half-bounded query ranges, into a single index, which eliminates the need to explicitly construct all $n$ HNSW indices. This is achieved by incrementally inserting items into an HNSW index in attribute-value order. Each edge in the resulting graph stores the attribute value at which it was added and, if pruned during construction, the attribute value at which it was removed. As a result, the HNSW graph contains edges that are valid only within specific attribute value intervals. When executing a query on the segment graph with a half-bounded query range, only edges whose validity interval contains the upper limit of the query range are considered. To generalize this approach for arbitrary query ranges (i.e., with both upper and lower limits), SeRF introduces the 2D segment graph (see Figure 10). This structure compresses $n$ segment graphs, one for each possible lower limit, into a single graph with $n$ vertices, where edges store two validity intervals: one for the lower and one for the upper limit of the query range. Additionally, SeRF proposes an in-filtering approach to extend its method to queries involving two or more attributes.
@@
iRangeGraph (Xu et al., 2024) (see Figure 11) is a graph-based FANNS index designed for range filtering. It is primarily built around a single ordered attribute but also supports in-filtering or post-filtering for secondary attributes of arbitrary types. iRangeGraph constructs a set of graph-based ANNS indices, referred to as elemental graphs. These elemental graphs are organized in a segment tree (De Berg, 2000) with $O(\log n)$ layers. At layer $l$, the entire attribute-value range is divided into $2^{l}$ segments, and a graph-based ANNS index is built for each segment, ensuring that each item appears once in each of the $\log n$ layers. Query execution employs an efficient on-the-fly method to dynamically merge a subset of elemental graphs into a single graph-based ANNS index containing only items within the query range. While its index construction is similar to $\beta$-WST (Engels et al., 2024), iRangeGraph differs significantly in query execution: instead of querying multiple graph-based indices for different subsets of items and merging the results, it constructs a single graph-based ANNS index dynamically at query time.
Figure 11. 3.3.17) Visualization of the iRangeGraph index.
@@ 3.3.18. UNIFY
UNIFY (Liang et al., 2024) is a graph-based method that supports range filtering. It introduces two new types of proximity graph, the segmented inclusive graph (SIG) and its HNSW-like variant the hierarchical segmented inclusive graph (HSIG) (see Figure 12). UNIFY assumes a single, ordered attribute. It divides the range of possible attribute values into $s$ segments, s.t., each segment contains approximately the same number of items. The high-level idea of a SIG is to build a proximity graph (Wang et al., 2021b) for each of the $2^{s}-1$ possible combinations of segments, and then merge these graphs into a single graph that contains the $2^{s}-1$ smaller graphs as subgraph. Due to a clever construction algorithm, the SIG can be built without explicitly constructing all $2^{s}-1$ smaller graphs. The SIG is stored as a segmented adjacency list, where outgoing edges are grouped by their destination segment. When querying the SIG, we only traverse the subgraph that correspond to those segments that intersect with the query range. The HSIG is a multi-layered variant of the SIG which is traversed from top to bottom in order to speed-up the query execution. To handle queries with low selectivity, UNIFY supports a pre-filtering variant and for queries with high selectivity, it supports unfiltered ANNS with post-filtering.
Figure 12. 3.3.18) UNIFY’s HSIG index visualized.
@@ 3.3.19. UNG
Figure 13. 3.3.19) Visualization of the UNG index.
@@ 4. A New Dataset to Benchmark FANNS
Benchmarking FANNS methods requires suitable datasets. Such datasets should contain vectors from a state-of-the-art embedding model (requirement 1), real-world attributes (requirement 2), a set of queries (vectors and filters) together with the pre-computed ground truth (requirement 3), and be publicly available (requirement 4).
@@
The most common approach to evaluate FANNS methods that we observed in scientific literature is to use an ANNS dataset such as SIFT (Laurent Amsaleg, 2025), GIST (Laurent Amsaleg, 2025), MNIST (LeCun, 2025), GloVe (Pennington et al., 2025), UQ-V (Fu, 2025), Msong (Fu, 2025), Audio (Fu, 2025), Crawl (Fu, 2025), Enron (Fu, 2025), BIGANN (Simhadri et al., 2025), Paper (Fu, 2025), Deep1B (Dmitry Baranchuk, 2025), MSTuring (Simhadri et al., 2025), or YandexT2I (Simhadri et al., 2025) and extend it with synthetic, randomly generated attributes (Xu et al., 2020; Wang et al., 2021d; Wang et al., 2024; Yang et al., 2020c; Zhao et al., 2022; Mohoney et al., 2023; Gollapudi et al., 2023; Gupta et al., 2023; Engels et al., 2024; Zuo et al., 2024; Liang et al., 2024; Cai et al., 2024) (failing to meet requirement 2). Another common, though more labor-intensive, approach is to repurpose a dataset not originally intended for FANNS such as TripClick (Rekabsaz et al., 2025), LAION (Schuhmann, 2025), RedCaps (Desai et al., 2025), YouTube-Rgb (Abu-El-Haija et al., 2025), Youtube-Audio (Abu-El-Haija et al., 2025), Words (Farrall, 2025), MTG (Trevor, 2025), or WIT-Image (Datasets, 2025) by manually creating embedding vectors and queries, or by crawling the web to obtain suitable real-world attributes (Patel et al., 2024; Engels et al., 2024; Zuo et al., 2024; Xu et al., 2024; Liang et al., 2024; Cai et al., 2024) (failing to meet requirement 3). Finally, some papers originating from industry rely on proprietary in-house datasets that are not publicly available (Wei et al., 2020; Yang et al., 2020c; Mohoney et al., 2023) (violating requirement 4).
The only two datasets that appear to satisfy all four requirements are MNIST (LeCun, 2025) and Paper (Fu, 2025). The MNIST dataset, where the labels correspond to the digits 0 to 999, and the Paper dataset, where each attribute has at most three possible values, cover only a narrow subset of potential FANNS use cases. Moreover, we observe that datasets with text embeddings from state-of-the-art, transformer-based models such as NV-Embed (Lee et al., 2024; NVIDIA, 2025), LENS (Lei et al., 2025; Lei, 2025), GTE (Li et al., 2023; Alibaba-NLP, 2025), Stella (Zhang et al., 2024; NovaSearch, 2025), and BGE (Chen et al., 2024; of Artificial Intelligence, 2025) are absent from the current ANNS and FANNS benchmarks (not in line with requirement 1). These models typically generate normalized embedding vectors with around 4096 dimensions, representing a fundamentally different FANNS workload compared to existing datasets, which usually contain embeddings with fewer than 1000 dimensions. We therefore argue that an additional datasets with transformer-based embeddings would be highly valuable to the community.

### v2

@@ 3.1. Filtering Approaches
Figure 2. 3.1) The best approach for solving FANNS depends on the filter’s selectivity within a given dataset.
The efficiency of different approaches to solving the FANNS problem depends on the dataset and filtering condition (see Figure 2). We define the selectivity of a filter $f$ on a dataset $D$ as the fraction of items in $D$ that satisfy the filter $f$, following previous work [^1] Note that some works (Wei et al., 2020; Wang et al., 2021d; Xu et al., 2024) use the inverse definition of selectivity. (Mohoney et al., 2023; Zhang et al., 2023; Patel et al., 2024; Liang et al., 2024). Pre-filtering is most efficient when selectivity is low, meaning only a few items remain for the ENNS. Post-filtering and in-filtering are most efficient when selectivity is high, i.e., most items pass the filter. If selectivity is too low, ANNS may not return enough valid candidates for post-filtering, and graph traversal with in-filtering may fail due to a sparsely connected or disconnected graph. A hybrid index is most efficient when selectivity is moderate, making none of the three previous approaches optimal.
3.2. Filter Types
We identify three fundamental filter types based on a single attribute. In Exact match (EM) filtering, the item’s attribute value must equal a specified query value. While EM filtering is usually applied to a categorical attribute, it can also be used on numerical attributes. Range (R) filtering applies to numerical attributes and considers only items where the attribute value is within a specified query range. Finally, Exact match in set (EMIS) filtering applies to set-valued attributes and considers only items whose attribute set contains a queried value. FANNS algorithms that apply the same filter type to multiple attributes implement multiple exact match (MEM), multiple range (MR), and multiple exact match in set (MEMIS) filters. We define a combined (C) filter as a filter that can combine the three fundamental single-attribute filters using arbitrary logical operators and may also support arbitrary SQL-style filter predicates.
3.3. Overview of FANNS Methods
In Table 1, we organize existing FANNS methods according to our taxonomy. We observe that most FANNS methods are based on graphs or quantization, with trees often used in combination with quantization- or graph-based techniques. Figure 3 groups FANNS methods according to their internal working principles, and in the following, we describe these methods in more detail.
Figure 3. 3.3) Grouping of FANNS method based on their internal working principles.
@@ 3.3.5. Milvus
Milvus (Wang et al., 2021d) extends the FAISS ANNS library (Douze et al., 2024; Research, 2025) by incorporating range filtering for one or multiple attributes. It supports both pre- and post-filtering, as well as a filtering scheme that partitions the data based on the most frequently filtered attribute and builds a separate ANNS index for each partition.
@@ 3.3.6. NHQ
NHQ (Wang et al., 2024) proposes a method to extend any proximity graph-based ANNS index into a hybrid FANNS index. Instead of constructing the ANNS index solely based on embedding distance, it introduces a fusion distance that combines embedding distance with attribute dissimilarity. While this approach is applicable to any proximity graph-based ANNS index, NHQ also presents construction schemes for two novel proximity graphs. Since its query execution scheme operates using the fusion distance without an explicit filtering step, NHQ may return items that do not match the filter, thus addressing the AFANNS rather than the FANNS problem.
@@ 3.3.7. HQANN
Like NHQ (Wang et al., 2024), HQANN (Yang et al., 2020c) employs a fusion distance to transform any proximity graph-based ANNS index into a hybrid FANNS index. The key difference to NHQ is that HQANN prioritizes attribute dissimilarity in the fusion distance, whereas NHQ emphasizes embedding distance.
@@ 3.3.8. AIRSHIP
AIRSHIP (Zhao et al., 2022) (see Figure 4) is a graph-based FANNS index capable of handling arbitrary filter functions. It relies on a graph-based ANNS index constructed without considering attributes but adapts query execution to FANNS through in-filtering and a series of optimizations. During query execution, AIRSHIP traverses the entire graph but includes only vertices that satisfy the filter in the results. It selects a starting point within a cluster of items that match the filter and traverses the graph in multiple directions concurrently.
@@ 3.3.9. HQI
Apple’s HQI (Mohoney et al., 2023) (see Figure 5) is a hybrid FANNS index. To integrate attribute filtering with nearest neighbor search, HQI transforms embedding vectors into attributes by applying k-means clustering and storing each item’s cluster ID as an attribute. It then employs a balanced qd-tree (Yang et al., 2020a), an attribute-only index, to partition the items. Each non-leaf node in the qd-tree contains a predicate over the item attributes (including the cluster ID), with its two child nodes holding items that do or do not satisfy the predicate. During query execution, HQI identifies the query vector’s $w$ closest cluster IDs and prunes all branches of the tree that do not contain any of these $w$ cluster IDs or whose attributes do not match the filter. An IVF index is used within each leaf.
@@ 3.3.10. VBASE
VBASE (Zhang et al., 2023) extends PostgreSQL (Group, 2025) by integrating ANNS into the database system. Its key insight is that most ANNS query execution methods follow a two-stage process: an initial stage where traversal moves from a starting point in the direction of the query vector, and a second stage where the search gradually moves away from the query vector to identify its KNN. Rather than employing a naïve pre-filtering approach, where the ANNS index is queried with a fixed $k^{\prime}\geq k$ and non-matching items are filtered out, VBASE runs the ANNS query only until the traversal starts moving away from the query vector. From that point onward, it iteratively retrieves and filters additional items until the required $k$ matches are found.
@@ 3.3.11. FDANN
Microsoft’s Filtered-DiskANN (FDANN) (Gollapudi et al., 2023) (see Figure 6) implements a graph-based index with EMIS filtering. In addition to supporting an EMIS filter with a single query-label, it also allows filtering with multiple query-labels, where at least one of them must be present in the item’s set attribute. Filtered-DiskANN introduces two methods for constructing FANNS indices based on Vamana graphs (Jayaram Subramanya et al., 2019): FilteredVamana and StitchedVamana. During query execution, it traverses the graph starting from a set of entry points that contain the query-label. Throughout its NSW-like graph traversal, it considers only those vertices that satisfy the filter criteria.
@@ 3.3.12. CAPS
In contrast to most FANNS methods, CAPS (Gupta et al., 2023) (see Figure 7) introduces a quantization- and tree-based index that is easier to parallelize than the more common graph-based indices. The CAPS index consists of two levels. The first level organizes items into clusters based on their embedding distance only or a combination of embedding distance and attribute dissimilarity. At the second level, CAPS constructs an attribute frequency tree (AFT), which is an attribute-only index. Each level of the AFT splits items into two buckets: one containing items with the most common attribute value (a leaf node) and another with the remaining items (a non-leaf node), unless the maximum depth is reached. During query execution, CAPS selects the $w$ clusters closest to the query vector and uses the corresponding AFTs to identify matching items.
@@ 3.3.13. ARKGraph
Given a set of items, each with a single ordered attribute, ARKGraph (Zuo and Deng, 2023) (see Figure 8) addresses the more general problem of constructing the K-nearest-neighbor graph (KGraph) (Dong et al., 2011) for the subset of items that satisfy a range filter. The resulting KGraph can be utilized for solving the FANNS problem or for other data analysis tasks. A naïve approach would store $O(n^{2})$ KGraphs, one for each possible query range. For any given item $I$ with attribute value $x$, this would require maintaining $O(n^{2})$ different lists of KNN, one for each possible query range $[l,r]$ such that $l\leq x\leq r$. ARKGraph’s first key insight is that the KNN of $I$ for any query range $[l,r]$ can be efficiently computed by merging the KNN of $I$ from the partial query ranges $[l,x)$ and $(x,r]$. Thus, instead of storing KNN lists for all $O(n^{2})$ query ranges, it suffices to maintain KNN lists for only $O(n)$ partial query ranges and reconstruct the KNN dynamically for any given range. The second key insight is that an item often shares the same KNN across multiple similar partial query ranges. By grouping query ranges and storing only one KNN set per group, ARKGraph further reduces index size and storage overhead.
@@ 3.3.14. ACORN
ACORN (Patel et al., 2024) (see Figure 9) supports arbitrary filter types and filtering across multiple attributes. The ACORN index is a denser variant of the HNSW ANNS index (Malkov and Yashunin, 2018), constructed without considering attributes. During query execution, ACORN traverses only the subgraph induced by vertices that match the filter. This approach differs from AIRSHIP (Zhao et al., 2022), which traverses the entire graph but includes only matching vertices in the result set. The ACORN index’s subgraphs approximate an HNSW (Malkov and Yashunin, 2018).
@@ 3.3.15. $\beta$-WST
$\beta$-WST (Engels et al., 2024) (see Figure 10) is a FANNS index designed for range filtering on a single attribute. A $\beta$-WST index is structured as a segment tree (De Berg, 2000) with a uniform branching factor of $\beta$. At layer $l$, the attribute value range is divided into $\beta^{l}$ segments, with a graph-based ANNS index constructed for each segment. The default query strategy selects a minimal set of tree nodes that fully cover the query range, queries the corresponding ANNS indices, and merges the results. Additional strategies include OptimizedPostfiltering, TreeSplit, and SuperPostfiltering.
@@ 3.3.16. SeRF
SeRF (Zuo et al., 2024) introduces the segment graph, a data structure for range-filtered FANNS queries with half-bounded query ranges (i.e., where the query range has an upper limit but no lower limit). It losslessly compresses $n$ HNSW indices (Malkov and Yashunin, 2018), corresponding to the $n$ possible half-bounded query ranges, into a single index. This is achieved by incrementally inserting items into an HNSW index in attribute-value order. Each edge in the resulting graph stores the attribute value at which it was added and, if pruned during construction, the attribute value at which it was removed. As a result, the HNSW graph contains edges that are valid only within specific attribute value intervals. During query execution, only edges whose validity interval contains the upper limit of the query range are considered. For query ranges with an upper and a lower limit, SeRF introduces the 2D segment graph (see Figure 11). This structure compresses $n$ segment graphs, one for each possible lower limit, into a single graph with $n$ vertices, where edges store validity intervals for the upper and lower limit of the query range.
@@
3.3.17. TFANNS
TFANNS (Luo et al., 2025) proposes three graph-based FANNS indices for EMIS filtering: The local-, global-, and packing method.
@@ 3.3.17. TFANNS
iRangeGraph (Xu et al., 2024) (see Figure 12) is a graph-based FANNS index designed for range filtering. It is primarily built around a single ordered attribute but also supports in-filtering or post-filtering for secondary attributes of arbitrary types. iRangeGraph constructs a set of graph-based ANNS indices, referred to as elemental graphs. These elemental graphs are organized in a segment tree (De Berg, 2000) with $O(\log n)$ layers. At layer $l$, the entire attribute-value range is divided into $2^{l}$ segments, and a graph-based ANNS index is built for each segment, ensuring that each item appears once in each of the $\log n$ layers. Query execution employs an efficient on-the-fly method to dynamically merge a subset of elemental graphs into a single ANNS index containing only items within the query range. While its index construction is similar to $\beta$-WST (Engels et al., 2024), iRangeGraph differs in query execution: instead of querying multiple graph-based indices for different subsets of items and merging the results, it constructs a single graph-based ANNS index dynamically at query time.
Figure 12. 3.3.18) Visualization of the iRangeGraph index.
3.3.19. RangePQ
RangePQ (Zhang et al., 2025) answers R-filtered FANNS queries by combining a PQ index for embedding vectors with a binary search tree (BST) for attribute values.
@@ 3.3.20. UNIFY
UNIFY (Liang et al., 2024) is a graph-based method that supports range filtering for a single attribute. It introduces two new types of proximity graph, the segmented inclusive graph (SIG) and its HNSW-like variant the hierarchical segmented inclusive graph (HSIG) (see Figure 13). It divides the range of possible attribute values into $s$ segments, s.t., each segment contains approximately the same number of items. The high-level idea of a SIG is to build a proximity graph (Wang et al., 2021b) for each of the $2^{s}-1$ possible combinations of segments, and then merge these graphs into a single graph that contains the $2^{s}-1$ smaller graphs as subgraph. Due to a clever construction algorithm, the SIG can be built without explicitly constructing all $2^{s}-1$ smaller graphs. The SIG is stored as a segmented adjacency list, where outgoing edges are grouped by their destination segment. When querying the SIG, we only traverse the subgraph that correspond to those segments that intersect with the query range. The HSIG is a multi-layered variant of the SIG which is traversed from top to bottom in order to speed-up the query execution. To handle queries with low selectivity, UNIFY supports a pre-filtering variant and for queries with high selectivity, it supports unfiltered ANNS with post-filtering.
Figure 13. 3.3.20) UNIFY’s HSIG index visualized.
@@ 3.3.21. UNG
Figure 14. 3.3.21) Visualization of the UNG index.
3.3.22. DIGRA
DIGRA (Jiang et al., 2025) uses a multi-way tree structure to organize items based on the value of their single ordered attribute. At each node of this tree, a NSW (Malkov et al., 2014) graph is used to perform ANNS on its subtree. Each R filtered query can be answered by querying only two of these NSW graphs.
3.3.23. DSG
The Dynamic Segment Graph (DSG) (Peng et al., 2025) is an extension of SeRF (Zuo et al., 2024) that is optimized for cases, where the vectors are dynamically inserted into the index in random order, and first sorting them by attribute value is not possible.
@@ 4. A New Dataset to Benchmark FANNS
Benchmarking FANNS methods requires suitable datasets. Ideally, such datasets should contain real-world attributes (criterion 1), queries with ground truth (criterion 2), and be publicly available (criterion 3). The most common approach to evaluate FANNS methods that we observed in scientific literature is to use an ANNS dataset such as SIFT (Laurent Amsaleg, 2025), GIST (Laurent Amsaleg, 2025), GloVe (Pennington et al., 2025), UQ-V (Fu, 2025), Msong (Fu, 2025), Audio (Fu, 2025), Crawl (Fu, 2025), Enron (Fu, 2025), BIGANN (Simhadri et al., 2025), Deep1B (Dmitry Baranchuk, 2025), MSTuring (Simhadri et al., 2025), or YandexT2I (Simhadri et al., 2025) and extend it with synthetic, randomly generated attributes (Xu et al., 2020; Wang et al., 2021d, 2024; Yang et al., 2020c; Zhao et al., 2022; Mohoney et al., 2023; Gollapudi et al., 2023; Gupta et al., 2023; Engels et al., 2024; Zuo et al., 2024; Liang et al., 2024; Cai et al., 2024; Jiang et al., 2025; Zhang et al., 2025) (failing to meet criterion 1). Another common, though more labor-intensive, approach is to repurpose a dataset not originally intended for FANNS such as TripClick (Rekabsaz et al., 2025), MNIST (LeCun, 2025), LAION (Schuhmann, 2025), RedCaps (Desai et al., 2025), YouTube-Rgb (Abu-El-Haija et al., 2025), Youtube-Audio (Abu-El-Haija et al., 2025), Words (Farrall, 2025), MTG (Trevor, 2025), or WIT-Image (Datasets, 2025) by manually creating embedding vectors and queries, or by crawling the web to obtain suitable real-world attributes (Patel et al., 2024; Engels et al., 2024; Zuo et al., 2024; Xu et al., 2024; Liang et al., 2024; Cai et al., 2024; Jiang et al., 2025) (failing to meet criterion 2). Finally, some papers originating from industry rely on proprietary in-house datasets that are not publicly available (Wei et al., 2020; Yang et al., 2020c; Mohoney et al., 2023) (violating criterion 3). The only dataset we are aware of that meets all three criteria is the Paper dataset  (Fu, 2025), containing 200-dimensional embeddings of research paper abstracts.
In addition to the scarcity of suitable datasets that fulfill all three criteria, we observe that text embeddings from state-of-the-art transformer-based models such as NV-Embed (Lee et al., 2024; NVIDIA, 2025), LENS (Lei et al., 2025; Lei, 2025), GTE (Li et al., 2023; Alibaba-NLP, 2025), Stella (Zhang et al., 2024; NovaSearch, 2025), and BGE (Chen et al., 2024; of Artificial Intelligence, 2025) with thousands of dimensions are not represented in existing FANNS datasets. To clarify whether such a dataset is desperately needed or if the performance of FANNS methods is agnostic to the embedding model used, we need to answer the following two questions:
  1. (1)
    Do the characteristics of embedding vectors vary significantly across different embedding models?
  2. (2)
    Does the performance of FANNS methods depend on the characteristics of the embedding vectors?
To answer the first question, we compare embeddings from the SIFT (Laurent Amsaleg, 2025) (images), Audio (Fu, 2025) (audio), UQ-V (Fu, 2025) (video), and Paper (Fu, 2025) (text) datasets to transformer-based text embeddings. In Table 2 we report the dimensionality of embedding vectors, the mean absolute correlation between dimensions of the embedding space, as well as the mean, standard deviation, and best-fitting distribution (among ten candidates) for vector norms and pairwise distances between vectors in each dataset . Figures 15, 16 and 17 show histograms for the vector norms and pairwise distances, as well as the correlation matrix for inter-dimension correlations of all five datasets. These results reveal our first key observation:
Observation [Uncaptioned image]
The characteristics of embedding vectors vary significantly across different embedding models.
Table 2. 4) Summary statistics of embedding vectors.
Dataset Quantity Mean ($\pm$ Std) Best Fit
SIFT Dimensions 128
Correlation 0.182
Vector Norms 509 $\pm$ 0.66 Beta (a=5.5,b=8.7a{=}5.5,b{=}8.7)
Pairw. Distances 536 $\pm$ 115 Beta (a=168,b=2.1a{=}168,b{=}2.1)
Audio Dimensions 192
Correlation 0.155
Vector Norms $7.9{\times}10^{5}\,\pm\,2.0{\times}10^{4}$ Beta (a=12.2,b=10.1a{=}12.2,b{=}10.1)
Pairw. Distances $1.46{\times}10^{5}\,\pm\,3.8{\times}10^{4}$ Normal
UQ-V Dimensions 256
Correlation 0.738
Vector Norms 0.225 $\pm$ 0.267 Lognormal (s=0.81s{=}0.81)
Pairw. Distances 0.201 $\pm$ 0.201 Lognormal (s=0.71s{=}0.71)
Paper Dimensions 200
Correlation 0.192
Vector Norms 2.13 $\pm$ 0.14 Lognormal (s=0.31s{=}0.31)
Pairw. Distances 1.10 $\pm$ 0.24 Beta (a=472,b=19.6a{=}472,b{=}19.6)
Ours (trans- former- based) Dimensions 4096
Correlation 0.099
Vector Norms 1.00 $\pm$ $1.3{\times}10^{-7}$ Normal
Pairw. Distances 1.08 $\pm$ 0.14 Weibull ($k{=}4.6{\times}10^{4}$)
Figure 15. 4) Distribution of vector norm among sampled embedding vectors.
Figure 16. 4) Distribution of pairwise distances among sampled embedding vectors.
Figure 17. 4) Correlation between different dimensions among sampled embedding vectors.
Figure 18. 4) Recall@10 vs. QPS plots for EM filtering on four widely used datasets of different modalities.
To answer the second question, we evaluate seven representative FANNS methods on the SIFT (Laurent Amsaleg, 2025), Audio (Fu, 2025), UQ-V (Fu, 2025), and Paper (Fu, 2025) datasets. For these experiments, we follow the same methodology as described in Section 5. Figure 18 shows the recall vs. QPS trade-off for each method on each dataset and Figure 19 shows the corresponding index construction time, peak memory usage, and index size. We observe that while some methods such as UNG or FDANN (both stitched and filtered) perform consistently across datasets, many methods exhibit significant performance variations depending on the dataset used. ACORN is significantly slower on Paper compared to other datasets, NHQ (nsw) fails to achieve a recall of over 0.4 on the Audio dataset, NHQ (kgraph) achieves significantly higher recall on SIFT and Paper than on Audio and UQ-V, and CAPS is way more competitive on Paper and Audio than on SIFT and UQ-V, which leads us to our second observation:
Observation [Uncaptioned image]
The performance of most FANNS methods depends on the characteristics of the embedding vectors.
Our first two observations motivate the need for a new FANNS dataset that includes embeddings from state-of-the-art transformer-based models. To this end, we introduce the arxiv-for-fanns dataset, where each item corresponds to a research paper. Our dataset is available in three scales: small (1k items, 197 MB), medium (100k items, 1.86 GB), and large (over 2.7M items, 46.1 GB). Each item includes a 4096-dimensional, normalized text embedding of the paper abstract, computed using the stella_en_400M_v5 model (NovaSearch, 2025; Zhang et al., 2024), based on abstracts from the arXiv dataset (Cornell University et al., 2025). We select stella because it is the most lightweight among the top 10 models [^2] As of March 11, 2025 on the English MTEB leaderboard (Muennighoff et al., 2022). Critically, each item contains 11 real-world attributes (see Table 3) covering categorical, numerical and set-valued attributes with diverse value distributions (shown in Figure 20), fulfilling criterion 1. To satisfy criterion 2, each dataset variant includes three sets of 10k queries for EM, R, and EMIS filters (see Section 3.2), along with their corresponding precomputed ground truth. Our queries filter for the number of sub categories (EM), update date (R), and for the main categories (EMIS), but a user can easily construct his own queries using any combination of the 11 attributes.